# Rates of convergence in long time asymptotics of an alignment model with symmetry breaking

### v2

@@ 5.2. A coercivity result
$\partial_{t}f=G_{u_{0}}\partial_{t}g^{0}=G_{u_{0}}(\mathcal{L}_{u_{0}}g^{0}+\mathcal{R}_{u_{0}}g^{0}).$
@@
$\partial_{t}f=u_{*}^{\prime}\nabla_{u_{*}}(G_{u_{*}})(1+g)+G_{*}\partial_{t}g.$
@@
$\nabla_{u}G_{u}(v)=\frac{v-u}{D}G_{u}(v)$ (44)
and $u_{*}^{\prime}\cdot u_{*}=0$. Since $u_{f}=u_{*}+\int_{\mathbb{R}^{d}}g\,v\,G_{*}\,dv$ and $u_{*}\in\mathcal{S}$, we obtain $u_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}g\,v\,G_{*}\,dv=0$. ∎
The advantage of considering the perturbation around $G_{u_{*}}$ instead of $G_{u_{\infty}}$ is that $v_{g}$ and $u_{*}$ are always colinear, hence Lemma˜5.3, Lemma˜5.4 and Lemma˜5.5 are rewritten as
$Q_{2,u_{*}}[g]\geq\beta(D)^{2}Q_{1,u_{*}}[g],$ (45)
$Q_{1,u_{*}}[g]\geq\eta(D)^{2}\lVert g\rVert_{2,u_{*}}^{2},$ (46)
@@
$a(D)Q_{2,u_{*}}[g]^{\frac{1}{2}}\leq\lVert\nabla g\rVert_{2,u_{*}}\leq b(D)Q_{2,u_{*}}[g]^{\frac{1}{2}}.$ (47)
@@
Under the assumptions of Theorem˜1.2, if $f$ solves (1) and $g$ and $u_{*}$ are defined by (42) and (30), then
$\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{u_{*}}=-Q_{2,u_{*}}[g]+R[g]$ (48)
@@
$R[g]$ $=D^{2}v_{g}\int_{\mathbb{R}^{d}}g(\nabla g-v_{g})G_{u_{*}}-\frac{u_{*}^{\prime}}{2}\cdot\int_{\mathbb{R}^{d}}g^{2}vG_{u_{*}}dv.$ (49)
@@
$\begin{split}\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{u_{*}}&=\langle g,\partial_{t}g\rangle_{u_{*}}+\frac{1}{2}\nabla_{u}\langle g,g\rangle_{u}\cdot u_{*}^{\prime}.\end{split}$
@@
$\begin{split}\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{u_{*}}&=-Q_{2,u_{*}}[g]+D^{2}\,v_{g}\cdot\int_{\mathbb{R}^{d}}g(\nabla g-v_{g})G_{*}\,dv\\ &\quad-\frac{1}{D}\langle v\cdot u_{*}^{\prime}(1+g),g\rangle_{u_{*}}+\frac{1}{2}\nabla_{u}\langle g,g\rangle_{u}\cdot u_{*}^{\prime}\end{split}$ (50)
Let us start with the computation of $\langle v\cdot u_{*}^{\prime}(1+g),g\rangle_{u_{*}}$. By definition (33),
$\langle v\cdot u_{*}^{\prime},g\rangle_{u_{*}}$ $=D\,u_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}v\,g\,G_{*}\,dv-D\int_{\mathbb{R}^{d}}(v\cdot u_{*}^{\prime})v\,G_{*}\,dv\cdot v_{g}.$
Assume $v_{g}\neq 0$ and $u_{*}^{\prime}\neq 0$. Since $v_{g}\cdot u_{*}^{\prime}=0$, we can consider an orthonormal basis $\{e_{1},e_{2},..,e_{d}\}$ with $e_{1}=u_{*}^{\prime}/|u_{*}^{\prime}|$ and $e_{2}=v_{g}/|v_{g}|$. In these coordinates, by writing $v=\sum_{i=1}^{d}v_{i}e_{i}$, we have $v\cdot u_{*}^{\prime}=v_{1}|u_{*}^{\prime}|$ and $v\cdot v_{g}=v_{2}|v_{g}|$. Since the function
$v_{1}\mapsto v_{1}\,v_{2}\,G_{*}\left(\sum_{i=1}^{d}v_{i}e_{i}\right)=v_{1}\,v_{2}\exp\left\{-\frac{1}{D}\left[\frac{\alpha}{4}\left(\sum_{i=1}^{d}|v_{i}|^{2}\right)^{2}-\frac{\alpha}{2}\left(\sum_{i=1}^{d}|v_{i}^{2}|\right)+\sum_{i=1}^{d}|v_{i}-u_{f}\cdot e_{i}|^{2}\right]\right\}$
is odd for all $(v_{2},v_{3},...v_{d})\in\mathbb{R}^{d-1}$ because $u_{f}\cdot e_{i}=0$ for $i\geq 2$, we obtain
$|v_{g}||u_{*}^{\prime}|\int_{\mathbb{R}^{d}}(v\cdot u_{*}^{\prime})v\,G_{*}\,dv\cdot v_{g}=\int\cdots\int\,dv_{2}\cdots dv_{d}\left(\int_{\mathbb{R}}v_{2}v_{1}\,G_{*}\left(\sum_{i=1}^{d}v_{i}e_{i}\right)dv_{1}\right)=0,$
thanks to the change of variable $(v_{1},v_{2},...v_{d})\mapsto\sum_{i=1}^{d}v_{i}e_{i}$. Since $u_{*}^{\prime}\cdot v_{g}=0$ we conclude
$\langle v\cdot u_{*}^{\prime},g\rangle_{u_{*}}=0.$
@@
$\langle v\cdot u_{*}^{\prime}(1+g),g\rangle_{u_{*}}=\langle v\cdot u_{*}^{\prime}g,g\rangle_{u_{*}}=D\,u_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}v\,g^{2}G_{*}\,dv-Dv_{g}\cdot\int_{\mathbb{R}^{d}}(v\cdot u_{*}^{\prime})\,v\,g\,G_{*}\,dv.$ (51)
Let us compute $\nabla_{u}\langle g,g\rangle_{u}$. Using (33) and the expression (44), we obtain:
$\nabla_{u}\langle g,g\rangle_{u}$ $=D\,\int_{\mathbb{R}^{d}}g^{2}\,\nabla_{u}G_{u}\,dv-\nabla_{u}\left|\int_{\mathbb{R}^{d}}g\,v\,G_{u}\,dv\right|^{2}$
$=\int_{\mathbb{R}^{d}}g^{2}(v-u)G_{u}+2\int_{\mathbb{R}^{d}}(v\otimes v)\,v_{g}\,g\,G_{u}\,dv-2\,D|v_{g}|^{2}u,$ (52)
@@
Finally, since $u_{*}^{\prime}\cdot u=0$, replacing (51) and (5.3) in (50),
$\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{u_{*}}$ $=-Q_{2,u_{*}}[g]+D^{2}v_{g}\cdot\int_{\mathbb{R}^{d}}g(\nabla g-v_{g})G_{{*}}\,dv$
$-\left[\,u_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}v\,g^{2}G_{*}\,dv-v_{g}\cdot\int_{\mathbb{R}^{d}}(v\cdot u_{*}^{\prime})\,v\,g\,G_{*}\,dv\right]$
$+\frac{1}{2}u_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}g^{2}\,v\,G_{*}\,dv+u_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}(v\otimes v)\,v_{g}\,g\,G_{*}\,dv.$
Using that $v_{g}\cdot(v\cdot u_{*}^{\prime})v=u_{*}^{\prime}\cdot(v\otimes v)v_{g}$, we obtain the result. ∎
@@
There exists a constant $C_{D}>0$ such that for all $u\in\mathcal{S}$ and $g\in H^{1}(\mathbb{R}^{d},G_{u}dv)$, there holds
$\frac{\alpha^{2}}{8D^{2}}\int_{\mathbb{R}^{d}}g^{2}|v|^{6}G_{u}dv\leq C_{D}\int_{\mathbb{R}^{d}}g^{2}G_{u}dv+\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{u}dv.$
@@
We adapt a classical method. Let $h=g\,G_{u}^{\frac{1}{2}}$. We recall that $G_{u}$ is given by (2). Let us denote  $\phi_{u}(v)=\frac{1}{D}(\frac{1}{2}|v-u|^{2}+\frac{\alpha}{4}|v|^{4}-\frac{\alpha}{2}|v|^{2})$}. We have,
$\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{u}dv$ $=\int_{\mathbb{R}^{d}}\left(|\nabla h|^{2}+\nabla h\cdot\nabla\phi_{u}h+\frac{1}{4}|\nabla\phi_{u}|^{2}h^{2}\right)dv$
$=\int_{\mathbb{R}^{d}}|\nabla h|^{2}dv+\int_{\mathbb{R}^{d}}h^{2}\left(\frac{1}{4}|\nabla\phi_{u}|^{2}-\frac{1}{2}\Delta\phi_{u}\right)dv.$
@@
$|\nabla\phi_{u}|^{2}=\frac{\alpha^{2}}{D^{2}}|v|^{6}+P_{u}(v)\text{ and }\Delta\phi_{u}(v)=Q_{u}(v)$
with $P_{u}$ and $Q_{u}$ respectively two polynomials (with $d$ variables) of degree at most than $5$ and $2$. Hence we can write
$\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{u}dv\geq\frac{\alpha^{2}}{4D^{2}}\int_{\mathbb{R}^{d}}g^{2}|v|^{6}G_{u}dv+\int_{\mathbb{R}^{d}}g^{2}S_{u}(v)G_{u}dv$
where $S_{u}(v)=\frac{1}{4}P_{u}(v)-\frac{1}{2}Q_{u}(v)$ is a polynomial of degree at most than $5$.
Let us take $R=R_{D}$ big enough in order that, for all $u\in\mathcal{S}$ and all $|v|\geq R$, there holds
$|S_{u}(v)|\leq\frac{\alpha^{2}}{8D^{2}}|v|^{6}.$
@@
$\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{u}dv$ $\geq\frac{\alpha^{2}}{8D^{2}}\int_{\mathbb{R}^{d}}g^{2}|v|^{6}G_{u}dv-\sup_{v\in B_{R}}S_{u}(v)\int_{\mathbb{R}^{d}}g^{2}G_{u}dv.$
Here $B_{R}$ denotes the ball of radius $R$ centered in 0. Hence we obtain the result with the constant  $C_{D}=\sup_{u\in\mathcal{S}}\sup_{v\in B_{R_{D}}}|S_{u}(v)|$. ∎
@@
Under the assumptions of Theorem˜1.2. Assume that for all $t\geq 0$, we have $u_{f}(t)\in\mathcal{N}$, then
@@
We have that $u_{*}(t)=\pi(u_{f}(t))$ where $\pi(v):=r(D)\frac{v}{|v|}$ is the projection on $\mathcal{S}$. Using the expression (29),
$u_{*}^{\prime}(t)$ $=d\pi_{u_{f}(t)}u_{f}^{\prime}(t)=d\pi_{u_{f}(t)}\left(-\alpha u_{f}(t)+\alpha\int_{\mathbb{R}^{d}}|v|^{2}vfdv\right)=d\pi_{u_{f}(t)}\left(\alpha\int_{\mathbb{R}^{d}}|v|^{2}vgG_{*}dv\right),$
where $d\pi_{v}(w)=\frac{w}{\|v\|}-\frac{v\cdot w}{\|v\|^{3}}v$ denotes the differential of $\pi$. Since $\lVert d\pi_{u_{f}}\rVert\leq\frac{1}{|u_{f}|}\leq\frac{2}{r(D)}$, we obtain
$|u_{*}^{\prime}(t)|\leq\frac{2\alpha}{r(D)}\int_{\mathbb{R}^{d}}|v|^{3}gG_{*}dv\leq\frac{2\alpha}{r(D)}\left(\int_{\mathbb{R}^{d}}|v|^{6}G_{*}dv\right)^{\frac{1}{2}}\lVert g\rVert_{2,u_{*}}$ $\leq\frac{2\alpha}{r(D)\eta(D)}W_{6}^{\frac{1}{2}}Q_{1,u_{*}}[g]^{\frac{1}{2}},$
@@
Moreover, if $u_{f}(t)\in\mathcal{N}$ for all $t\geq 0$,
$R[g(t,.)]\leq K_{D}Q_{2,u_{*}}[g(t,.)]Q_{1,u_{*}}^{\frac{1}{2}}[g(t,.)]\,\,\forall t\geq 0,$ (54)
@@
$R[g]$ $=D^{2}v_{g}\int_{\mathbb{R}^{d}}g(\nabla g-v_{g})G_{*}-\frac{u_{*}^{\prime}}{2}\cdot\int_{\mathbb{R}^{d}}g^{2}\,v\,G_{*}\,dv.$
Let us compute $v_{g}\cdot\int_{\mathbb{R}^{d}}\,g\,(\nabla g-v_{g})\,G_{{*}}$. Using (45) and (46),
$\left|D^{2}v_{g}\cdot\int_{\mathbb{R}^{d}}g(\nabla g-v_{g})G_{{*}}\right|$ $\leq D^{2}|v_{g}|\lVert g\rVert_{2,u_{*}}\lVert\nabla g-v_{g}\rVert_{2,u_{*}}$
$\leq\frac{D^{2}}{\eta(D)\beta(D)}|v_{g}|Q_{2,u_{*}}[g].$
Let us compute $\int_{\mathbb{R}^{d}}g^{2}vG_{u_{*}}dv.$
$\left|\int_{\mathbb{R}^{d}}g^{2}vG_{{*}}dv\right|$ $\leq\left(\int_{\mathbb{R}^{d}}g^{2}|v|^{6}G_{{*}}\right)^{\frac{1}{6}}\left(\int_{\mathbb{R}^{d}}g^{2}G_{{*}}\right)^{\frac{5}{6}}$
@@
$\left|\frac{u_{*}^{\prime}}{2}\cdot\int_{\mathbb{R}^{d}}g^{2}vG_{*}dv\right|\leq\frac{\gamma(D)}{2}|u_{*}^{\prime}|Q_{2,u_{*}}[g].$
@@
Equation˜54 follows from (53), using Lemma˜5.9 and the inequality
$|v_{g}|\leq\frac{1}{D\eta(D)}W_{2}^{1/2}Q_{1,u_{*}}[g]^{1/2}.\qed$
@@
Under the assumptions of Theorem˜1.2 there exist some constants  $A_{D}>0$ and  $K_{D}>0$ such that if $Q_{1,u_{*}(0)}[g_{\mathrm{ini}}]<A_{D}$, then
${Q_{1,u_{*}}[g]}\leq\frac{Q_{1,u_{*}}[g_{\mathrm{ini}}]}{\left(1-K_{D}Q_{1,u_{*}}[g_{\mathrm{ini}}]^{\frac{1}{2}}\right)^{2}}e^{-2\beta(D)^{2}t}\,\,\,\forall t\geq 0$
@@
$|u_{f}(t)-u_{*}(t)|\leq\delta\,\,\forall t\geq 0.$

### v3

@@ 5.2. A coercivity result
$\partial_{t}f=G_{\mathbf{u}_{0}}\partial_{t}g^{0}=G_{\mathbf{u}_{0}}(\mathcal{L}_{\mathbf{u}_{0}}g^{0}+\mathcal{R}_{\mathbf{u}_{0}}g^{0}).$
@@
$\partial_{t}f=\mathbf{u}_{*}^{\prime}\nabla_{\mathbf{u}_{*}}(G_{\mathbf{u}_{*}})(1+g)+G_{*}\partial_{t}g.$
@@
$\nabla_{\mathbf{u}}G_{\mathbf{u}}(v)=\frac{v-\mathbf{u}}{D}G_{\mathbf{u}}(v)$ (45)
and $\mathbf{u}_{*}^{\prime}\cdot\mathbf{u}_{*}=0$. Since $\mathbf{u}_{f}=\mathbf{u}_{*}+\int_{\mathbb{R}^{d}}g\,v\,G_{*}\,dv$ and $\mathbf{u}_{*}\in\mathcal{S}$, we obtain $\mathbf{u}_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}g\,v\,G_{*}\,dv=0$. ∎
The advantage of considering the perturbation around $G_{\mathbf{u}_{*}}$ instead of $G_{\mathbf{u}_{\infty}}$ is that $\mathbf{v}_{g}$ and $\mathbf{u}_{*}$ are always colinear, hence Lemma˜5.3, Lemma˜5.4 and Lemma˜5.5 are rewritten as
$Q_{2,\mathbf{u}_{*}}[g]\geq\beta(D)^{2}Q_{1,\mathbf{u}_{*}}[g],$ (46)
$Q_{1,\mathbf{u}_{*}}[g]\geq\eta(D)^{2}\lVert g\rVert_{2,\mathbf{u}_{*}}^{2},$ (47)
@@
$a(D)Q_{2,\mathbf{u}_{*}}[g]^{\frac{1}{2}}\leq\lVert\nabla g\rVert_{2,\mathbf{u}_{*}}\leq b(D)Q_{2,\mathbf{u}_{*}}[g]^{\frac{1}{2}}.$ (48)
@@
Under the assumptions of Theorem˜1.2, if $f$ solves (1) and $g$ and $\mathbf{u}_{*}$ are defined by (43) and (31), then
$\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{\mathbf{u}_{*}}=-Q_{2,\mathbf{u}_{*}}[g]+R[g]$ (49)
@@
$R[g]$ $=D^{2}\mathbf{v}_{g}\int_{\mathbb{R}^{d}}g(\nabla g-\mathbf{v}_{g})G_{\mathbf{u}_{*}}\,dv-\frac{\mathbf{u}_{*}^{\prime}}{2}\cdot\int_{\mathbb{R}^{d}}g^{2}vG_{\mathbf{u}_{*}}dv.$ (50)
@@
$\begin{split}\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{\mathbf{u}_{*}}&=\langle g,\partial_{t}g\rangle_{\mathbf{u}_{*}}+\frac{1}{2}\nabla_{\mathbf{u}}\langle g,g\rangle_{\mathbf{u}}\cdot\mathbf{u}_{*}^{\prime}.\end{split}$
@@
$\begin{split}\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{\mathbf{u}_{*}}&=-Q_{2,\mathbf{u}_{*}}[g]+D^{2}\,\mathbf{v}_{g}\cdot\int_{\mathbb{R}^{d}}g(\nabla g-\mathbf{v}_{g})G_{*}\,dv\\ &\quad-\frac{1}{D}\langle v\cdot\mathbf{u}_{*}^{\prime}(1+g),g\rangle_{\mathbf{u}_{*}}+\frac{1}{2}\nabla_{\mathbf{u}}\langle g,g\rangle_{\mathbf{u}}\cdot\mathbf{u}_{*}^{\prime}\end{split}$ (51)
Let us start with the computation of $\langle v\cdot\mathbf{u}_{*}^{\prime}(1+g),g\rangle_{\mathbf{u}_{*}}$. By definition (34),
$\langle v\cdot\mathbf{u}_{*}^{\prime},g\rangle_{\mathbf{u}_{*}}$ $=D\,\mathbf{u}_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}v\,g\,G_{*}\,dv-D\int_{\mathbb{R}^{d}}(v\cdot\mathbf{u}_{*}^{\prime})v\,G_{*}\,dv\cdot\mathbf{v}_{g}.$
Assume $\mathbf{v}_{g}\neq 0$ and $\mathbf{u}_{*}^{\prime}\neq 0$. Since $\mathbf{v}_{g}\cdot\mathbf{u}_{*}^{\prime}=0$, we can consider an orthonormal basis $\{e_{1},e_{2},..,e_{d}\}$ with $e_{1}=\mathbf{u}_{*}^{\prime}/|\mathbf{u}_{*}^{\prime}|$ and $e_{2}=\mathbf{v}_{g}/|\mathbf{v}_{g}|$. In these coordinates, by writing $v=\sum_{i=1}^{d}v_{i}e_{i}$, we have $v\cdot\mathbf{u}_{*}^{\prime}=v_{1}|\mathbf{u}_{*}^{\prime}|$ and $v\cdot\mathbf{v}_{g}=v_{2}|\mathbf{v}_{g}|$. Since the function
$v_{1}\mapsto v_{1}\,v_{2}\,G_{*}\left(\sum_{i=1}^{d}v_{i}e_{i}\right)=v_{1}\,v_{2}\exp\left\{-\frac{1}{D}\left[\frac{\alpha}{4}\left(\sum_{i=1}^{d}|v_{i}|^{2}\right)^{2}-\frac{\alpha}{2}\left(\sum_{i=1}^{d}|v_{i}^{2}|\right)+\sum_{i=1}^{d}|v_{i}-\mathbf{u}_{f}\cdot e_{i}|^{2}\right]\right\}$
is odd for all $(v_{2},v_{3},...v_{d})\in\mathbb{R}^{d-1}$ because $\mathbf{u}_{f}\cdot e_{i}=0$ for $i\geq 2$, we obtain
$|\mathbf{v}_{g}||\mathbf{u}_{*}^{\prime}|\int_{\mathbb{R}^{d}}(v\cdot\mathbf{u}_{*}^{\prime})v\,G_{*}\,dv\cdot\mathbf{v}_{g}=\int\cdots\int\,dv_{2}\cdots dv_{d}\left(\int_{\mathbb{R}}v_{2}v_{1}\,G_{*}\left(\sum_{i=1}^{d}v_{i}e_{i}\right)dv_{1}\right)=0,$
thanks to the change of variable $(v_{1},v_{2},...v_{d})\mapsto\sum_{i=1}^{d}v_{i}e_{i}$. Since $\mathbf{u}_{*}^{\prime}\cdot\mathbf{v}_{g}=0$ we conclude
$\langle v\cdot\mathbf{u}_{*}^{\prime},g\rangle_{\mathbf{u}_{*}}=0.$
@@
$\langle v\cdot\mathbf{u}_{*}^{\prime}(1+g),g\rangle_{\mathbf{u}_{*}}=\langle v\cdot\mathbf{u}_{*}^{\prime}g,g\rangle_{\mathbf{u}_{*}}=D\,\mathbf{u}_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}v\,g^{2}G_{*}\,dv-D\mathbf{v}_{g}\cdot\int_{\mathbb{R}^{d}}(v\cdot\mathbf{u}_{*}^{\prime})\,v\,g\,G_{*}\,dv.$ (52)
Let us compute $\nabla_{\mathbf{u}}\langle g,g\rangle_{\mathbf{u}}$. Using (34) and the expression (45), we obtain:
$\nabla_{\mathbf{u}}\langle g,g\rangle_{\mathbf{u}}$ $=D\,\int_{\mathbb{R}^{d}}g^{2}\,\nabla_{\mathbf{u}}G_{\mathbf{u}}\,dv-\nabla_{\mathbf{u}}\left|\int_{\mathbb{R}^{d}}g\,v\,G_{\mathbf{u}}\,dv\right|^{2}$
$=\int_{\mathbb{R}^{d}}g^{2}(v-\mathbf{u})G_{\mathbf{u}}\,dv+2\int_{\mathbb{R}^{d}}(v\otimes v)\,\mathbf{v}_{g}\,g\,G_{\mathbf{u}}\,dv-2\,D|\mathbf{v}_{g}|^{2}\mathbf{u},$ (53)
@@
Finally, since $\mathbf{u}_{*}^{\prime}\cdot\mathbf{u}=0$, replacing (52) and (5.3) in (51),
$\frac{1}{2}\frac{d}{dt}\langle g,g\rangle_{\mathbf{u}_{*}}$ $=-Q_{2,\mathbf{u}_{*}}[g]+D^{2}\mathbf{v}_{g}\cdot\int_{\mathbb{R}^{d}}g(\nabla g-\mathbf{v}_{g})G_{{*}}\,dv$
$-\left[\,\mathbf{u}_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}v\,g^{2}G_{*}\,dv-\mathbf{v}_{g}\cdot\int_{\mathbb{R}^{d}}(v\cdot\mathbf{u}_{*}^{\prime})\,v\,g\,G_{*}\,dv\right]$
$+\frac{1}{2}\mathbf{u}_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}g^{2}\,v\,G_{*}\,dv+\mathbf{u}_{*}^{\prime}\cdot\int_{\mathbb{R}^{d}}(v\otimes v)\,\mathbf{v}_{g}\,g\,G_{*}\,dv.$
Using that $\mathbf{v}_{g}\cdot(v\cdot\mathbf{u}_{*}^{\prime})v=\mathbf{u}_{*}^{\prime}\cdot(v\otimes v)\mathbf{v}_{g}$, we obtain the result. ∎
@@
There exists a constant $C_{D}>0$ such that for all $\mathbf{u}\in\mathcal{S}$ and $g\in H^{1}(\mathbb{R}^{d},G_{\mathbf{u}}dv)$, there holds
$\frac{\alpha^{2}}{8D^{2}}\int_{\mathbb{R}^{d}}g^{2}|v|^{6}G_{\mathbf{u}}dv\leq C_{D}\int_{\mathbb{R}^{d}}g^{2}G_{\mathbf{u}}dv+\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{\mathbf{u}}dv.$
@@
Let $h=g\,G_{\mathbf{u}}^{\frac{1}{2}}$. We recall that $G_{\mathbf{u}}$ is given by (2). Let us denote $\phi_{\mathbf{u}}(v)=\frac{1}{D}(\frac{1}{2}|v-\mathbf{u}|^{2}+\frac{\alpha}{4}|v|^{4}-\frac{\alpha}{2}|v|^{2})$}. We have,
$\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{\mathbf{u}}dv$ $=\int_{\mathbb{R}^{d}}\left(|\nabla h|^{2}+\nabla h\cdot\nabla\phi_{\mathbf{u}}h+\frac{1}{4}|\nabla\phi_{\mathbf{u}}|^{2}h^{2}\right)dv$
$=\int_{\mathbb{R}^{d}}|\nabla h|^{2}dv+\int_{\mathbb{R}^{d}}h^{2}\left(\frac{1}{4}|\nabla\phi_{\mathbf{u}}|^{2}-\frac{1}{2}\Delta\phi_{\mathbf{u}}\right)dv.$
@@
$|\nabla\phi_{\mathbf{u}}|^{2}=\frac{\alpha^{2}}{D^{2}}|v|^{6}+P_{\mathbf{u}}(v)\text{ and }\Delta\phi_{\mathbf{u}}(v)=Q_{\mathbf{u}}(v)$
with $P_{\mathbf{u}}$ and $Q_{\mathbf{u}}$ respectively two polynomials (with $d$ variables) of degree at most than $5$ and $2$. Hence we can write
$\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{\mathbf{u}}dv\geq\frac{\alpha^{2}}{4D^{2}}\int_{\mathbb{R}^{d}}g^{2}|v|^{6}G_{\mathbf{u}}dv+\int_{\mathbb{R}^{d}}g^{2}S_{\mathbf{u}}(v)G_{\mathbf{u}}dv$
where $S_{\mathbf{u}}(v)=\frac{1}{4}P_{\mathbf{u}}(v)-\frac{1}{2}Q_{\mathbf{u}}(v)$ is a polynomial of degree at most than $5$.
Let us take $R=R_{D}$ big enough in order that, for all $\mathbf{u}\in\mathcal{S}$ and all $|v|\geq R$, there holds
$|S_{\mathbf{u}}(v)|\leq\frac{\alpha^{2}}{8D^{2}}|v|^{6}.$
@@
$\int_{\mathbb{R}^{d}}|\nabla g|^{2}G_{\mathbf{u}}dv$ $\geq\frac{\alpha^{2}}{8D^{2}}\int_{\mathbb{R}^{d}}g^{2}|v|^{6}G_{\mathbf{u}}dv-\sup_{v\in B_{R}}S_{\mathbf{u}}(v)\int_{\mathbb{R}^{d}}g^{2}G_{\mathbf{u}}dv.$
Here $B_{R}$ denotes the ball of radius $R$ centered in 0. Hence we obtain the result with the constant $C_{D}=\sup_{\mathbf{u}\in\mathcal{S}}\sup_{v\in B_{R_{D}}}|S_{\mathbf{u}}(v)|$.
@@
Under the assumptions of Theorem˜1.2. Assume that for all $t\geq 0$, we have $\mathbf{u}_{f}(t)\in\mathcal{N}$ , where $\mathcal{N}$ is defined by (22), then
@@
We have that $\mathbf{u}_{*}(t)=\pi(\mathbf{u}_{f}(t))$ where $\pi(v):=r(D)\frac{v}{|v|}$ is the projection on $\mathcal{S}$. Using the expression (30),
$\mathbf{u}_{*}^{\prime}(t)$ $=d\pi_{\mathbf{u}_{f}(t)}\mathbf{u}_{f}^{\prime}(t)=d\pi_{\mathbf{u}_{f}(t)}\left(-\alpha\mathbf{u}_{f}(t)+\alpha\int_{\mathbb{R}^{d}}|v|^{2}vfdv\right)=d\pi_{\mathbf{u}_{f}(t)}\left(\alpha\int_{\mathbb{R}^{d}}|v|^{2}vgG_{*}dv\right),$
where $d\pi_{v}(w)=\frac{w}{|v|}-\frac{v\cdot w}{|v|^{3}}v$ denotes the differential of $\pi$. Since $\lVert d\pi_{\mathbf{u}_{f}}\rVert\leq\frac{1}{|\mathbf{u}_{f}|}\leq\frac{2}{r(D)}$, we obtain
$|\mathbf{u}_{*}^{\prime}(t)|\leq\frac{2\alpha}{r(D)}\int_{\mathbb{R}^{d}}|v|^{3}gG_{*}dv\leq\frac{2\alpha}{r(D)}\left(\int_{\mathbb{R}^{d}}|v|^{6}G_{*}dv\right)^{\frac{1}{2}}\lVert g\rVert_{2,\mathbf{u}_{*}}$ $\leq\frac{2\alpha}{r(D)\eta(D)}W_{6}^{\frac{1}{2}}Q_{1,\mathbf{u}_{*}}[g]^{\frac{1}{2}},$
@@
Moreover, if $\mathbf{u}_{f}(t)\in\mathcal{N}$ for all $t\geq 0$,
$R[g(t,.)]\leq K_{D}Q_{2,\mathbf{u}_{*}}[g(t,.)]Q_{1,\mathbf{u}_{*}}^{\frac{1}{2}}[g(t,.)]\,\,\forall t\geq 0,$ (55)
@@
$R[g]$ $=D^{2}\mathbf{v}_{g}\cdot\int_{\mathbb{R}^{d}}g(\nabla g-\mathbf{v}_{g})G_{*}\,dv-\frac{\mathbf{u}_{*}^{\prime}}{2}\cdot\int_{\mathbb{R}^{d}}g^{2}\,v\,G_{*}\,dv.$
Let us compute $\mathbf{v}_{g}\cdot\int_{\mathbb{R}^{d}}\,g\,(\nabla g-\mathbf{v}_{g})\,G_{{*}}$. Using (46) and (47),
$\left|D^{2}\mathbf{v}_{g}\cdot\int_{\mathbb{R}^{d}}g(\nabla g-\mathbf{v}_{g})G_{{*}}\,dv\right|$ $\leq D^{2}|\mathbf{v}_{g}|\lVert g\rVert_{2,\mathbf{u}_{*}}\lVert\nabla g-\mathbf{v}_{g}\rVert_{2,\mathbf{u}_{*}}$
$\leq\frac{D^{2}}{\eta(D)\beta(D)}|\mathbf{v}_{g}|Q_{2,\mathbf{u}_{*}}[g].$
Let us compute $\int_{\mathbb{R}^{d}}g^{2}vG_{\mathbf{u}_{*}}dv.$
$\left|\int_{\mathbb{R}^{d}}g^{2}vG_{{*}}dv\right|$ $\leq\left(\int_{\mathbb{R}^{d}}g^{2}|v|^{6}G_{{*}}\,dv\right)^{\frac{1}{6}}\left(\int_{\mathbb{R}^{d}}g^{2}G_{{*}}\,dv\right)^{\frac{5}{6}}$
@@
$\left|\frac{\mathbf{u}_{*}^{\prime}}{2}\cdot\int_{\mathbb{R}^{d}}g^{2}vG_{*}dv\right|\leq\frac{\gamma(D)}{2}|\mathbf{u}_{*}^{\prime}|Q_{2,\mathbf{u}_{*}}[g].$
@@
Equation˜55 follows from (54), using Lemma˜5.9 and the inequality
$|\mathbf{v}_{g}|\leq\frac{1}{D\eta(D)}W_{2}^{1/2}Q_{1,\mathbf{u}_{*}}[g]^{1/2}.\qed$
@@
Under the assumptions of Theorem˜1.2 there exist some constants $A_{D}>0$ and $K_{D}>0$ such that if $Q_{1,\mathbf{u}_{*}(0)}[g_{\mathrm{in}}]<A_{D}$, then
${Q_{1,\mathbf{u}_{*}}[g]}\leq\frac{Q_{1,\mathbf{u}_{*}}[g_{\mathrm{in}}]}{\left(1-K_{D}Q_{1,\mathbf{u}_{*}}[g_{\mathrm{in}}]^{\frac{1}{2}}\right)^{2}}e^{-2\beta(D)^{2}t}\,\,\,\forall t\geq 0$
@@
$|\mathbf{u}_{f}(t)-\mathbf{u}_{*}(t)|\leq\delta\,\,\forall t\geq 0.$